| Andrea Cunial | andrea.cunial@studenti.unipd.it | 2012-12-21 | 15 30 UT | 07 00 UT | 80% | Cirrus | High humidity at the start of the night and clouds deny observations for 2 hours more or less.
After that, no other problems. |
| Andrea Cunial | andrea.cunial@studenti.unipd.it | 2012-12-19 | 15 40 UT | 06 40 UT | 100% | Clear | No problems with the instrumentation.
Problems with the new weather station, that failed after the 21 UT. No information about external humidity was given. |
| Andrea Cunial | andrea.cunial@studenti.unipd.it | 2012-12-20 | 18 00 UT | 02 00 UT | 0% | Rain | Snow all the night.
New Weather Station was going all the night. |
| Andrea Cunial | andrea.cunial@studenti.unipd.it | 2012-12-22 | 17 00 UT | 01 30 UT | 0 % | Scattered Clouds | The telescope dome stopped at 22:30 UT at 41.2° azimut angle (direction South-West for the dome aperture). This problem denied the following observation. |
| Fiaschi-Angeloni | marco@mfc-elettronica.it | 2012-12-28 | 16:00 | 24 00 | 100 | Clear | |
| Fiaschi - Angeloni | marco@mfc-elettronica.it | 2012-12-29 | 16.00 | 23.30 | 100% | Clear | very good seeing |
| Fiaschi - Angeloni | marco@mfc-elettronica.it | 2012-12-30 | 16.00 | 22.30 | 100 | Clear | very clear until 22.30, variable seeing from 2 arcsec to 5 arcsec |
| Fiaschi - Angeloni | marco@mfc-elettronica.it | 2013-01-11 | 20.00 | 00.30 | 80% | Cirrus | |